OpenStudy (anonymous):

log_a(9) / 2log_a(27)

OpenStudy (anonymous):

ans it plzzzzzzzzzzzzzzzzzzzzzzzzzzz

OpenStudy (anonymous):

maybe use the change of base formula

OpenStudy (anonymous):

plz write how to do

OpenStudy (anonymous):

= loga9 / loga9^3 - i assume u want to simplify??

OpenStudy (anonymous):

yeah simplify

OpenStudy (anonymous):

the thing thats bothersome is the 2 in front of the log_a27

OpenStudy (anonymous):



OpenStudy (anonymous):

i think thats what it is

OpenStudy (anonymous):

let y = loga 9 9 = a^y z = loga729 = loga 9^3 log a^9 / loga^9^3 = 1/3 thats it 1/3

OpenStudy (anonymous):

@jimmy- yea thats what i got too. if u simplify log9/2log27, i got .3333333 aka 1/3

OpenStudy (anonymous):

right

OpenStudy (anonymous):

log_a(9)/2log_a(27) = log_a(3^2) / 2log_a(3^3) = 2 log_a(3) / 6 log_a(3) - Using the power identity = 2/6 = 1/3 Read Wikipedia: http://en.wikipedia.org/wiki/Logarithms#Product.2C_quotient.2C_power.2C_and_root